GBT Operations & Commissioning Meeting 05 December 2003 AGENDA 1. Az Track and Telescope Status -- Bob 2. Observing News -- Ron 3. Frontends, IFs, and Spectral Baselines -- Roger 4. Spectrometer Status -- Rich 5. Software Status -- Nicole 6. Scheduling Status -- Carl 7. Project Planning Status -- John 8. Any other Business PRELIMINARY REPORTS 1. Az Track and Telescope Status Azimuth Track * SGH continues to make progress on the models in jumps. They have the model of a wear plate on a baseplate up to 18" in just 27 hours of run time. Still having trouble with the other model - the software company has taken this on fully - SGH cannot make any progress on it. Model simply quits, with no error messages. Software company is going through it line by line, iteration by iteration to find the failure. * Tim Weadon, Bill Shank, and Greg Monk worked all day yesterday to collect the moving track measurements. This wraps up the vertical relative movement measurments of each track component. Another half day is needed for Harry Morton to obtain the absolute movements, and for Tim to measure relative movements from a tangential position. This will be scheduled for either next week or the week after, depending on other telescope work. * Requested an indication of availability and an hourly rate to perform the second UT of the modified joint. No response yet. Structural inspection * No word on Pete Keating's report from M&M. -- RAA 2. Observing News The following is the report that covers the last two weeks. Unfortunately, it's not been a good two weeks. I count that the telescope was to be used for 265 hours for either observing or tests. About 55 of those hours were lost to high winds. Of the remaining 210 hours, about 14 hours were lost to about 30 software or hardware failures. In previous weeks, both the time and number of failures was about three times lower than this. I'd have to go back to early September before I found a time when failures cost us so much time. o According to the operator's logs, over a dozen times either vortex, aio1, aio2, eio1, .... had to be restarted. On at least two occasions, we lost significant fractions of the system. o It took two hours to determine why LO1 was entering a "Fault" condition. The problem ended up being mostly insufficient error checking and not very helpful error messages. o Spectral Processor had to be reset about eight times. Spectrometer twice. o The antenna gave us PF boom or SafeHold problems about 4 times. o GlishD/Iards failed about 4 times. o Four times Naiad/Titania either logged people out or had their screen savers lock up. Other items to note: o We had a problem with a filter in the L-band receiver that 'fixed' itself. o Q-band commissioning was cut short by a dead channel. o K-band re-re-commissioning started again with so far glowing grades. -- RJM 3. Frontends, IFs, and Spectral Baselines 40-52 GHz Receiver: Initial results on the GBT were fairly encouraging. One channel was found to have serious refrigerator modulation, and the receiver was brought back to the lab. A tiny bit of metallic debris was located in the LNA input waveguide, and removed. The receiver was ready on Tuesday the 2nd to be reinstalled, but weather and other maintenance activities prevented installation this week. 18-26.5 GHz Receiver: The receiver was reinstalled on November 25, after a broken attenuator was replaced. 26.5-40 GHz Receiver: The internal dewar components have been assembled with test loads in place of the feed horns and the dewar has been vacuum tested with a cover in place of the feed window assembly. The dewar holds vacuum well, leaking less than 10 microns Hg in 24 hours, cooling and RF tests are planned within the next two weeks. Materials for a vacuum feed window have been obtained and will be tested for leaking and RF suitability in the next few weeks. The 290K section of the differencing engine has been assembled and is being tested for gain versus frequency response. Multiplexer assembly and tuning are proceeding, we have four multiplexers ready and are preparing two spares. Video amplifier and filter design is proceeding, a prototype is currently being assembled. (G. Watts) PF1: Pulled off on December 3 for cryo maintenance and repair of the filterbank in one channel. Other: The reworked feed defroster is ready to be installed when the antenna schedule allows. Work continues on constructing bandpass filters for the mmConverter, construction of additional LO Reference modules, and on the controller for the outdoor antenna range upgrade. -- RDN 12/4/2003 4. Spectrometer Status ****Status****: The spectrometer continues to operate fairly reliably. ****Development****: Spigot testing: Modes 15 and 16 have been confirmed to work by David Kaplan. These are the first non-1024 lag modes to be qualified. LTA replacement: Design of the FPGAs continues. Cross-correlation test fixture: The noise source circuit has been designed and simulated. A printed circuit is being designed for it. High Speed Samplers and Baselines: New cables have been fabricated for one sampler. The sampler passes its normal tests with the new cables. ****Plans****: Continue work on LTA replacement, cross-correlation test fixture and spigot mode testing. Modify one high speed sampler with better supply decoupling. (This should make the system ready for more baseline tests.) -- Rich and Holly 5. Software Status Single Dish Development IPT #61 - Friday, December 5, 2003 Today ends week 3 of 6 in the current development cycle, which is the 9th and last in 2003. The Plan of Record for the current development cycle is now available from the wiki at Software.PlanOfRecordC92003 along with the most up-to-date status information. A key accomplishment this week is that all build errors have been removed from the M&C system that are not the result of third party software. At the beginning of the year, there were over 2000 errors. Now, only two remain in the third party software, and these will be remedied when those modules are ported to Linux. Before the recent holiday, Paul and Eric successfully supported a PTCS demonstration of pointing and focus temperature corrections using a data display system that will eventually become the replacement for IARDS. Over the next few days, the requirements for making the corrections a production capability by the end of the cycle will be revisited based on the initial results. Work continues on defining an API for observing, which will culminate with at least two live tests by the end of the cycle (one for a continuum observation and one for a spectral line observation). Based on Frank and Jim's analysis, Mark and Amy are completing the design this week, which will keep us on schedule for proof of concept tests by the end of the cycle. Analysis is also underway for porting the converter rack and analog filter rack to Linux. Most members of the division attended the PTCS review this past Wednesday. Joe also provided one of the panel members with sample data sets and examples of antenna manager commands for review. Operational support this week included a fix to the LO1, which was behaving abnormally when a modified VLBI script was being used to run zero length scans (in an attempt to skip those scans). Response was also provided when Jim had difficulties with security in the beta version of the configuration tool. The production quality security mechanism has not been developed yet, but this must be in place prior to the production release of configuration, after which the problem will not occur. -- NMR 6. Scheduling Status Last Two Weeks ============= Observations for: GBT02C-058, GBT03A-002, GBT03C-031, GBT03B-026, GBT03B-030, GBT03A-023, GBT03B-015, GBT03A-016, BU027, GBT03C-023, GBT03C-014, GBT02C-020 Completed: GBT03C-023 December ======== Scheduled hours [backup]: Astronomy ~ 291 [181] Maintenance ~ 121 Tests & Comm ~ 283 [46] Shutdown ~ 36 Un assigned ~ 14 January ====== Scheduled hours [backup]: Astronomy ~ 389 [131] Maintenance ~ 9 [124] Tests & Comm ~ 203 Un assigned ~ 143 February ======= Scheduled hours [backup]: Astronomy ~ 196 Maintenance ~ [140] Tests & Comm ~ 115 Un assigned ~ 386 Proposal Checkouts =============== Awaiting checkout - 0 Active checkout - 30 Schedule - 46 On Hold - 5 Completed - 122 -- RCB 7. Project Planning Status December 1st Planning Meeting Minutes Rm 137 10:00 A.M. [0] Observer comments & operator log items Weather related lost time We discussed the GBT wind stow algorithm and the contribution to lost time this past week. We need to revisit this to be sure that we are not unnecessarily shutting down the telescope. Software related lost time Lost time on Holography run Converter rack lockups Spectrometer problems Spectral processor problems These items were discussed, and the are symptomatic of general system connectivity problems. SDD will be working on these problems as an operations response this week. [1] Resource calendar schedule conflicts and discussions None. Azimuth track measurements this week. Nothing scheduled for next week. 3 maintenance days per week for the next 3 weeks. [2] Observing Schedule discussions Unassigned time remains, as does some unassigned test time. Ron and Richard will meet and divvy up any unneeded PTCS test time. [3] GBT development planning 8-10 GHz receiver upgrade (Heiles & Robishaw's proposal) There is a proposal to add a polarization switch and a different filter to the X band receiver. This is to accomodate some measurements that are proposed by Heiles and Robishaw. The modifications will take some 3 weeks of engineering and 3 weeks of software maintenance to implement. We take this up at our 6 month planning meeting in March, and will attempt to fit it in this Spring. CPSR2 pulsar coherent dedispersion machine There is a proposal by Bryan Jacoby to bring a pulsar backend here. We will provide the IF, clock, and a network connection. Bryan will do the rest of the work. Phil will communicate with both of these groups and request a short written proposal so that we can be sure all of the bases are covered. [4] AOB None. -- JF 8. Any other Business